CLMult01: Reference Card Version 1.0

This is a list of commands provided by the LaTEX style files CLMult01. For a complete description of these commands please refer to the user’s guide.

Using the style file

tex &clmult01 <filename>

Documentstyle

clmult01

Standard Styleoptions

array, draft, makeidx, twoside, leqno, fleqn

Springer Styleoptions

amssymb, bibay, deutsch, footinfo, francais, german, numart, referee,
vecphys

Front Matter

The front matter for the volume is done by the editor(s) and contains the table of contents and the list of contributors.

\tableofcontents

List of Contributors

\begin{contributorlist}
{\bf <author name>}
<address>
\and
...
\end{contributorlist}

Title and Author

\title{<your title>}
\subtitle{<subtitle>}
\author{<author's name>}\inst{<number>}
\institute{<name of the first institute>
\and <name of the second institute> ...}
\maketitle

Changing Running Heads

\authorrunning{<name of first author> et al.}
\titlerunning{<short version of title>}

Must appear before the command \maketitle.

Lists with Variable Indentation

\begin{description}[<largelabel>]
\item[<label1>] <text1>
\item[<label2>] <text2>
...
\end{description}

Small Print

\begin{petit} <text> \end{petit}

Special Fonts and Symbols

\D   \E   \I
\vec{<symbol>}       \tens{<symbol>}    \frak{<symbol>}
{\cal <symbol>}      {\sf <symbol>}     {\Bbb <symbol>}

AMS symbols: see AMS-TEX guide

Mini-Environments Such as Lemmas and Theorems

\begin{<name of environment>}[<optional material>]
<text for that environment>
\end{<name of environment>}

Defining New Lemma-like Environments

\newtheorem{<env_nam>} {<caption>} [<within>] {<cap_font>}
  {<body_font>}
\newtheorem{<env_nam>} [<num_like>] {<caption>} {<cap_font>}
  {<body_font>}
\newtheorem*{<env_nam>}{<caption>}{<cap_font>}{<body_font>}

Figures

\begin{figure}
\sidecaption  % this is optional
\mpicplace{<width> cm}{<height> cm}
\caption{<text of caption>}
\end{figure}

Tables Not Coded with LaTEX

\begin{table}
\caption{<text of caption>}
\picplace{<height> cm}
\end{table}

Reference List (with bibay option)

\begin{thebibliography}[<label>]{<title>}
\bibitem{}{<key>}{<citation>} <bibliographic entry>
...
\end{thebibliography}

Index

\index{<entry>}
\index{<mainentry>!<subentry>}
\index{<mainentry>!<subentry>!<subsubentry>}

Sorting the Index (requires the MakeIndex program, the command is to be given at the operating system level)

makeindx -s sprmindx.sty <filename>

Printing the Index

\printindex

Manipulating Index Entries

change sorting order: \index{<text1>@<text2>}
other font: \index{major entry@{\it major entry}}
range: \index{<entry>|(} ... \index{<entry>|)}
macros inside pagenumber: \index{<entry>|<macrocall without \ >}

“see”: \see{<second entry>}

Semimanual Creation of an Index

Create file <filename>.idx containing lines

\indexentry{<keyword>}{<pageno>}

Manual Creation of an Index

\printindex in the document

Create dataset called <filename>.ind containing

\begin{theindex}
\item <entry>\idxquad <pagenumber(s)>
\subitem <entry>\idxquad <pagenumber(s)>
\subsubitem <entry>\idxquad <pagenumber(s)>
\indexspace
\end{theindex}

Appendices

\appendix

Acknowledgements

\begin{acknowledgement} <text> \end{acknowledgement}

Springer-Verlag

Berlin Heidelberg NewYork London Paris Tokyo Hong Kong Barcelona Budapest